# A deep graph model for the signed interaction prediction in biological network

### v1

@@ 2.3 Conflicting Edge Sampling Strategy
After constructing and evaluating multiple candidate models, we selected the best-performing model for further performance assessment. Our goal was to verify the model’s capability in predicting interactions between chemicals and genes and to understand the impact of different network connectivity on prediction accuracy. The best model’s ROC and PRC curves are shown in the figures (Fig. 2 A, D). In the Receiver Operating Characteristic (ROC) curves, the three lines illustrate the relationship between the true positive rate and the false positive rate at various thresholds. All three lines rapidly ascend to high true positive levels, indicating the model’s high sensitivity to positive cases. Especially, Curve 3 shows the steepest ascent before an increase in false positive rate, indicating its superior discriminative ability. The Precision-Recall (PRC) curves describe the relationship between precision and recall. Here, the closer a curve is to the upper right corner of the graph, the better the model’s performance. All curves maintain high precision at high recall rates, indicating the model’s ability to maintain high precision while keeping a high recall rate.
Additionally, we conducted tests on models employing the CL sampling strategy and those not utilizing any sampling strategy (hereafter referred to as the CL model and the No CL model, respectively) to observe their ability to identify different relational dynamics, particularly in distinguishing edges of opposite polarity. Using a test set of the same data volume, we predicted the probability of four types of relationships (Increase, Decrease, Binding, Affect) for each edge within the model. Notably, Increase and Decrease were considered relations of opposite polarity; the emergence of high probabilities simultaneously in both directions indicated a polarity conflict error, suggesting the model’s inability to differentiate edge polarity in the chemical-gene interactions.
Figure 2: Model test result: A,B show the ROC and PRC curves for CL sampling model. C and D provide the degree distribution of chemicals and genes, respectively. E shows the distribution of polarity degrees C CL model and NO CL model which is calculated by Equation(12) after mathematical transformation. F shows the contributions to polarity identify of both CL model and NO CL model, red and blue represent correctly and incorrectly prediction,separately.
For models before and after the implementation of the CL sampling strategy, we defined the Polarity Degree (C) as the absolute difference between the predicted probabilities of the ”Increase” and ”Decrease” directions for the same edge. We calculated the C values for both models and showed the frequency distribution of C values using histograms (Fig.2 E) and use Equation(12) to make transformation. It was observed that the model without the CL sampling strategy (represented in red) predominantly had a lower polarity degree, whereas the CL model (in blue) showed a higher polarity degree. This indicates that the CL model was more effective in managing or reducing contradictions and could more distinctly differentiate edge polarity.
Further, we marked correctly identified polar relationships with a +1 and incorrect or unrecognizable polar relationships with a -1 during the testing phase, depicted in red and blue lines, respectively. Through aggregate analysis of each line in the testing process (Fig.2 F), it was evident that in the CL model, occurrences and the extent of blue lines were significantly less frequent, indicating that instances of wrong predictions or errors were substantially fewer compared to the No CL model.
To validate the significant effect of our model on handling polar edges after employing the CL sampling strategy, we conducted a validation prediction on the model. We tested the model without the CL sampling strategy and with the CL sampling strategy on a randomly selected set of compounds, ranking the probability of each gene node across all connection scenarios. We separately tallied the top 100 nodes with the highest rankings in the polar relations of ”Increase” and ”Decrease”, converted the ranking values to log2, and plotted the log2-transformed rankings for each gene for increase and decrease scenarios (Fig.3 A-D). We sampled 900 compounds for repeated testing and plotted a boxplot of the average value of the log2 difference of the rankings of the top 100 nodes for each compound (Fig.3 E).
Figure 3: Polar Edge Analysis and CL Sampling Strategy Validation: A-D represent the log2-transformed ranking values of the top 100 nodes under conditions of increasing and decreasing interactions, with and without the CL sampling strategy applied. Panel E shows a boxplot of the average log2 rank differences in rankings of these nodes, based on repeated testing across 100 compounds.
From the figures, it’s evident that in the absence of CL sampling (NO_CL), both the top 100 rankings for increase and decrease show little difference in log2 rank values, and overlap exists. However, after applying CL sampling, the overlap significantly reduces and disappears. Looking at the slope boxplots obtained through repeated testing, after applying CL sampling (CL), the slope values of the boxplots significantly increase, indicating that the ranking changes between the ”Increase” and ”Decrease” states of gene activity are more pronounced. The increase in slope values, especially the higher absolute values observed in the CL boxplot, clearly demonstrates that CL sampling enhances the differentiation between data points, making the distinction clearer in the direction of chemical-gene activity. Overall, CL sampling significantly enhances the differentiation effect of chemical-gene interaction data, which is crucial for understanding and interpreting gene regulatory mechanisms.
@@ 2.4 Impact of Network Component
As detailed in previous sections, our prediction model and its unique features, including the sampling method and decoder structure we adopted, have been discussed. Specifically, when we attempted to introduce subgraph structures (only chemical-chemical subgraph, only gene-gene subgraph, and both subgraphs) for training to explore their potential impact on the predictive performance of chemical-gene data, the results could have met our expectations. We discovered that, in some instances, adding subgraph structures could actually have a detrimental effect on model performance. Specifically, without using subgraphs, our model’s AUROC, AUPRC, and AP@20 evaluation metrics scored 0.985, 0.985, and 1.00000, respectively. And our two new mertics shown 0.695 and 0.121. However, upon introducing subgraph structures, these metrics experienced a decline (Table 4), indicating that in our model, subgraph structures are not a beneficial addition for enhancing predictive performance and might instead lead to performance degradation due to overcomplexity. This finding poses important considerations for model design: even if specific approaches are effective in other models or datasets, they may not be universally applicable.
@@
Prediction $AUROC$ $AUPRC$ $AP@20$ $AUC_{polarity}$ $CP@500$
Subgraph 0.839 0.825 0.781 0.521 0.771
Only Chemical 0.948 0.938 0.836 0.546 0.972
Only Gene 0.917 0.882 0.697 0.557 0.810
No Subgraph 0.985 0.984 1.000 0.695 0.866
@@
This result highlights the independence and efficiency of our model in processing chemical-gene interaction data. It indicates that for the specific dataset, our model has been optimized to a point where introducing subgraph structures not only fails to enhance predictive performance but may even reduce it due to the introduction of unnecessary complexity. Our study underscores the importance of developing models tailored to a dataset’s specific needs and provides practical guidance on balancing increasing model complexity with maintaining or enhancing performance. Future work will be needed to further explore the relationship between model complexity and predictive performance, as well as how to adjust the model structure for chemical-gene datasets to maximize prediction accuracy.
@@ 3.1 Model Construction
Our model aims to compute the probability of the types of edges of interest given a specific node combination. This model employs an encoder-decoder architecture, where the encoder adopts the Graph Convolutional Network (GCN) principles designed by Yang et al. in the development of BioNet[23], and the decoder utilizes the design of Relational Graph Convolutional Networks (RGCN). During training, a strategy of sampling conflicting edges from signed and unsigned network combinations is adopted to address the diversity of edge types. Our model consists of two main components: GCNEncoder and RGCNTDDecoder. The GCNEncoder is divided into two layers; the first layer uses a graph convolutional layer to capture the features of the original input data; the second layer continues to abstract these features through another graph convolutional layer to form a more refined representation of the nodes. Subsequently, the RGCNTDDecoder combines Relational Graph Convolutional Network (RGCN) layers with the DEDICOM tensor decomposition layer, where the former is responsible for aggregating information according to different relation types, and the latter for parsing global and local interactions, ultimately generating predictions on the interactions between entities. The model’s architecture is shown in (Fig.4), which fully considers the multi-relational nature of graph data, enabling the capture of rich structural information and thereby holding extensive application potential in fields such as bioinformatics and recommendation systems.
@@ 3.1.1 Graph Convolutional Encoder
In this study, we employed a Graph Convolutional Encoder designed to capture and encode the intricate cooperative interactions between chemicals and genes. The graph convolutional encoder utilizes a hierarchical graph convolutional network architecture that iteratively aggregates and propagates information across nodes to learn node embeddings. The input to this architecture is a graph constructed from nodes represented as one-hot vectors and edges represented through adjacency matrices. In our work, we opt for a two-layer GCN optimized using Chebyshev polynomials, which has proven to exhibit enhanced performance, as evidenced in previous studies. The feature update for a node $v_{i}$ is given by Equation(1):
@@
Here, $h_{i}$ represents the final node embedding, with $h_{i}^{C}$ and $h_{i}^{G}$ respectively denoting the original embeddings of chemicals and genes. We merge these two types of node embeddings through a simple vector concatenation operation $\oplus$, instead of the multi-subgraph structure mentioned in the previous model. This design choice is based on the characteristics and needs of our data, aiming to simplify the model structure while preserving key information. Our model utilizes the hierarchical structure of the graph convolutional network during the encoding phase, first calculating the embeddings for each type of node separately and then combining these embeddings through the concatenation operation. Such a structure allows the model to reduce computational complexity without sacrificing performance, enabling us to train on larger-scale data.

### v2

@@ 2.3 Designed Metrics for Polairty Prediction
This result suggests that our custom metrics, AUC_polarity and CP@500, effectively evaluate model performance in polarity prediction scenarios. Not only do they confirm the effectiveness of the CL sampling strategy, but they also highlight the significant advantage of the RGCNTD model in addressing complex interactions and polarity edge prediction.
@@ 2.4 Impact of Network Component
Besides chemical-gene interaction networks, we also explored introducing subgraph structures (only chemical-chemical subgraphs, gene-gene subgraphs, and both) for training to understand their potential impact on the predictive performance of chemical-gene interactions. We discovered that, in some instances, adding subgraph structures could harm model performance (Table 4). Without the introduction of subgraph structures (No Subgraph), the model achieved its best performance across all metrics, with macro AUROC and micro AUROC reaching 0.985 and 0.980, respectively, macro AUPRC and micro AUPRC at 0.975 and 0.974, and AP@20 at 1. Additionally, the two custom metrics, AUC_polarity and CP@500, scored 0.696 and 0.866, respectively, demonstrating the model’s exceptional classification ability and prioritization performance. In contrast, when subgraph structures were introduced (Subgraph), the model’s performance significantly declined, with macro AUROC and micro AUROC dropping to 0.839 and 0.877, macro AUPRC and micro AUPRC reducing to 0.825 and 0.851, and AP@20 falling to 0.781. Furthermore, the metrics AUC_polarity and CP@500 also decreased to 0.521 and 0.771, respectively, indicating that the RGCNTD model is better suited for data structures without subgraphs.
@@
Prediction MACRO_AUROC MICRO_AUROC MACRO_AUPRC MICRO_AUPRC AP@20 $AUC_{\text{polarity}}$ CP@500
Subgraph 0.839 0.877 0.825 0.851 0.781 0.521 0.771
Only Chemical 0.948 0.933 0.938 0.927 0.836 0.546 0.972
Only Gene 0.917 0.910 0.882 0.891 0.697 0.557 0.810
No Subgraph 0.985 0.980 0.975 0.974 1 0.695 0.866
@@
This result highlights the independence and efficiency of the RGCNTD model in processing chemical-gene interaction data. It indicates that for the specific dataset, the RGCNTD model has been optimized to a point where introducing subgraph structures not only fails to enhance predictive performance but may even reduce it due to the introduction of unnecessary complexity. Our study underscores the importance of developing models tailored to a dataset’s specific needs and provides practical guidance on balancing increasing model complexity with maintaining or enhancing performance. Future work will be needed to explore further the relationship between model complexity and predictive performance and how to adjust the model structure for chemical-gene interaction datasets to maximize prediction accuracy.
@@ 3.1 Model Construction
The RGCNTD model is designed to compute the probability of edge types of interest based on specific node combinations. The model consists of two primary components: the GCN Encoder and the RGCN Tensor Decomposition Decoder. The GCN Encoder is divided into two layers; the first layer employs a graph convolutional layer to capture the features of the original input data, and the second layer further abstracts these features through another graph convolutional layer to form a more refined representation of the nodes. Subsequently, the RGCN Tensor Decomposition Decoder combines Relational Graph Convolutional Network (RGCN) layers with a DEDICOM tensor decomposition layer, where the former aggregates information according to different relationship types. At the same time, the latter parses global and local interactions, ultimately generating predictions for interactions between entities. As shown in (Fig.3), the model’s architecture fully accounts for the multi-relational nature of graph data, enabling the capture of rich structural information and thus holding extensive application potential in fields such as bioinformatics and recommendation systems.
@@
3.2 Network Construction
In this study, we constructed a network of interactions between chemicals and genes to train models using data from the STITCH database [22]. PubChem CIDs from STITCH were mapped to the CHEBI ID in the ChEBI database [23]. The ChEBI database excludes protein molecules encoded by genes and other insignificant small molecules, focusing solely on chemically meaningful entities. Moreover, ChEBI’s ontology provides a rich structure that facilitates understanding relationships between specific chemical entities and the broader chemical domain [24]. Similarly, protein IDs from STITCH were mapped to NCBI Gene ID.
We also introduced homogeneous graphs, such as the Chemical-Chemical and Gene-Gene interaction networks, to provide the model with additional information and evaluate the impact of complex network inputs on predictive performance. The Gene-Gene interaction network was derived from 1,307,492 interaction records in the STRING database[25]. Chemical-Chemical interaction data was obtained from 1,770,581 records in the CHEBI database[26]. Relationships were based on polarity into the following categories:
  1. [label=1]
  2. 1.
    Increase: The compound exerts an activation effect on gene expression or the activity of gene expression products.
  3. 2.
    Decrease: The compound exerts an inhibition effect on gene expression or the activity of gene expression products.
  4. 3.
    Binding: The compound binds to the gene expression pathway or the products of gene expression.
  5. 4.
    Affect: The compound affects gene expression or the activity of gene expression products, but the effect lacks clear polarity or its mechanism remains unclear.
After data cleaning and removing irrelevant molecules, the resulting heterogeneous Chemical-Gene graph contained 12,537 chemical nodes, 28,432 gene nodes, and 183,3943 interaction records distributed across the four relationship types (Table 5). Within this network, we analyzed pairs exhibiting multiple relationship types. We found that 103,738 pairs of entities have both polar and non-polar relationships.
Table 5: Relations in chemical-gene subgraph
No. Realtion Name Number of the edges Polarity
1 Increase 752,037 Increase Polarity
2 Decrease 667,187 Decrease Polarity
3 Binding 403,915 No Polarity
4 Affect 10,803 No Polarity
[…]
3.3 Handling Polar Edges in Chemical-Gene Prediction Tasks
In chemical-gene interaction prediction, our dataset encounters the challenge of handling polar edges. Polar edges refer to instances where a chemical and a gene (e.g., the relationship between e1 and e2) can exhibit contradictory effects (e.g., increased vs. decreased expression), typically considered mutually exclusive. To effectively capture and present these conflicting relationships, we designed a dual sampling strategy that incorporates both ‘Must-Link’ (ML) and ‘Cannot-Link’ (CL) constraints. These constraints are defined as Equation(1):
$\begin{split}\text{ML}=\left\{(e_{1},r_{\text{label}},e_{2}),(e_{1},r_{\text{predict}},e_{2})\mid r_{\text{label}}\in\text{Group}_{1},\,r_{\text{predict}}\in\text{Group}_{1}\right\}\\ \text{CL}=\left\{(e_{1},r_{\text{label}},e_{2}),(e_{1},r_{\text{predict}},e_{2})\mid r_{\text{label}}\in\text{Group}_{1},\,r_{\text{predict}}\in\text{Group}_{1}^{\prime}\right\}\end{split}$ (1)
Here, $r_{\text{label}}$ represents the true relationship label, and $r_{\text{predict}}$ represents the predicted relationship label by the model. $\text{Group}_{1}$ and $\text{Group}_{1}^{\prime}$ represent opposite types of relationships. For example, if $\text{Group}_{1}$ represents all relationships associated with ”increases” , then $\text{Group}_{1}^{\prime}$ represents the corresponding ”decreases” relationships. These definitions help the model more effectively manage similar or opposing relationship types during training. To reflect these constraints in model training, we designed the following loss function Equation(2):
$L=\sum_{\begin{subarray}{c}(e_{1},r_{\text{predict}},e_{2})\in\text{ML}\end{subarray}}\left(1-S(e_{1},r_{\text{predict}},e_{2})\right)^{2}+\sum_{\begin{subarray}{c}(e_{1},r_{\text{predict}},e_{2})\in\text{CL}\end{subarray}}\left(S(e_{1},r_{\text{predict}},e_{2})\right)^{2}$ (2)
In this loss function, the goal of the Must-Link (ML) constraint is to ensure that the score $S(e_{1},r_{\text{predict}},e_{2})$ for edges belonging to the same semantic category is close to 1. Conversely, the goal of the Cannot-Link (CL) constraint is to ensure that the score $S(e_{1},r_{\text{predict}},e_{2})$ for edges belonging to opposite semantic categories is close to 0. This design enables the model to accurately distinguish between different relationship types during training, effectively handling polar edges and improving the accuracy and reliability of the predictions.
In our implementation, we focused specifically on the polarity information of the ‘Increase’ and ‘Decrease’ relationship types, enhancing the model’s ability to differentiate mutually exclusive interactions. By integrating this mechanism into the training routine, the model could learn more nuanced data representations that capture the complexity of biological systems, where certain relationships are inherently polar. This approach improved the model’s discriminative power, resulting in more accurate and biologically meaningful predictions.
3.4 Baseline
We selected RGCN, GraphSAGE, and TransE as baseline models to compare their performance with our RGCNTD model on the benchmarks. RGCN utilizes relation-specific convolutions to handle multi-relational graphs, making it well-suited for predicting edge labels in multi-relational networks. GraphSAGE, known for its scalability, efficiently processes large graphs through neighbor sampling and feature aggregation. TransE, widely used for knowledge graph embeddings, models relationships via vector translations, providing a useful comparison for evaluating our model’s ability to handle complex edge-node interactions. The model parameters are summarized in Table 6:
Table 6: Model Parameters
No. Parameter GraphSAGE RGCN TransE
1 grad_norm 1.0 1.0 1.0
2 n_bases 4 4 4
3 learing rate 1e-2 1e-2 1e-2
4 regularization 1e-2 1e-2 1e-2
5 n_epochs 1000000 100000 1000000
[…]
3.5 Experimental Setup
All experiments were conducted on a server equipped with an NVIDIA GeForce RTX 3090. The RGCNTD model’s encoder was a two-layer graph convolutional network optimized with Chebyshev polynomials. The Adam optimizer was employed for optimization, with an early stopping mechanism incorporated to prevent overfitting. Early stopping was based on the validation loss, with a patience threshold of 5 epochs. The final model was trained for 50 epochs with a batch size of 8192, and the best model parameters were identified at epoch 48. A test size of 8192 was used, and the test results were recorded. A detailed list of parameters used in the model is provided in Table 7
Table 7: Model Parameters
No. Parameter Default Description
1 epoch 50 Number of epoches
2 hidden_dimensions [128, 64, 32, 16] List of hidden dimensions for layers
3 batch_size 8192 Batch size for training
4 test_batch True Whether to test the data per batch
5 test_batch_size 8192 Batch size for testing
6 print_step 10 Steps of printing information
7 average_precision_k 20 k for average precision at k
[…]
3.6 Metrics Design
In this study, we introduced a range of evaluation metrics to assess the models’ performance comprehensively, besides the Area Under the Receiver Operating Characteristic Curve (AUROC) and Area Under the Precision-Recall Curve (AUPRC) and AP@20. To account for the class imbalance in the samples, we calculated both macro and micro levels for the AUROC and AUPRC metrics. Additionally, we designed two novel metrics to evaluate the model’s ability to distinguish polar edges. These metrics highlight the model’s effectiveness in predicting chemical-gene interactions and offer valuable insights for future research.
3.6.1 Average Precision at Polar Edges
$C=|P_{\text{increase}}-P_{\text{decrease}}|$ (3)
Here, $P_{\text{increase}}$ and $P_{\text{decrease}}$ represent the model’s predicted probabilities of a chemical increasing or decreasing a gene’s activity, respectively. This metric quantifies the model’s confidence and capability in distinguishing between these two interaction types. However, the initial distribution of C values is typically long-tailed, reflecting a degree of uncertainty in the model’s predictions for these interactions. To improve the statistical properties of C and facilitate data analysis and interpretation, we applied a mathematical transformation. This transformation function is monotonically increasing within the range of 0 to 1, ensuring that the original ordering of C values is preserved without altering their relative magnitudes. To standardize the range of C and address its high dynamic range, we employed a natural logarithm function (ln). This approach compresses the dynamic range of C, resulting in a smoother and more uniform distribution. The transformation also incorporates a scaling constant, such as $2\pi^{2}$, which can be replaced with other alternatives if needed. The transformation function is specified as Equation(4):
$C^{\prime}=\frac{\ln(1+2\pi^{2}\cdot C)}{\ln(2\pi^{2}+1)}$ (4)
We observed significant changes in the distribution of C before and after the transformation. Before transformation(Fig.4 A), the distribution of C exhibited a long-tailed shape, particularly in the non-CL model. A dense concentration of low C values indicated substantial uncertainty in the model’s ability to distinguish between ’Increase’ and ’Decrease’ interactions. The long-tailed nature of the distribution makes visual inspection more difficult, as the frequent occurrence of extreme values obscures meaningful patterns in the data. After transformation (Fig.4 B), the distribution of C improved significantly, especially within the CL model. The transformed distribution showed reduced long-tail effects and display greater symmetry and concentration. This adjustment made the distribution easier to visualize and analyze while preserving the metric’s original properties.
3.6.2 Polarity Degree (C)-based Precision
To comprehensively assess the model’s performance in high-confidence predictions, we introduce the Polarity Degree (C)-based Precision at 500 (CP@500) metric. This metric evaluates the model’s accuracy among its top 500 most confident predictions, measuring its precision under high-confidence conditions. The calculation involves the following steps:
  1. [label=1]
  2. 1.
    1. For each instance in the dataset, we compute the polarity degree (C) as defined in Equation
  3. 2.
    The predictions are then sorted in descending order based on their C value, prioritizing those with the highest confidence (largest C values)
  4. 3.
    3. Each prediction is evaluated for correctness by comparing the predicted category (the category with larger predicted probability) to the true category. A prediction is marked as correct (1) if the predicted category matches the true category; otherwise, it is marked incorrect (0)
  5. 4.
    4. Finally, the top k predictions from the sorted list are selected, and the number of correct predictions among these is used to compute CP@k as shown in Equation(5):
[…]
By introducing CP@k, we not only evaluate the model’s accuracy in its most confident predictions, but also emphasize its performance under high-confidence conditions. This metric provides an intuitive and effective way to compare the precision of different models in their top-ranked predictions.
3.6.3 AUC for polar edge
Inspired by the AUC (Area Under the Curve) evaluation methodology, we devised a novel metric called Area Under the Curve for polar edge prediction (AUC_polarity), specifically designed to evaluate the model performance in predicting the polarity (’Increase’ or ’Decrease’) of relationships between nodes in a network. To compute AUC_polarity, we randomly select N pairs of nodes connected by edges labeled with ’Increase’ or ’Decrease’ polarity. The model predicts the probabilities of ’Increase’ $p_{a}$ and ’Decrease’ $p_{a}$, and the task is to determine the actual polarity of each edge based on these probabilities. We utilize a binary indicator $\hat{y}$. $\hat{y}$ is assigned a value of 1 if the model’s predicted probability of ’Increase’ is greater than that of ’Decrease’ and the actual polarity is ’Increase’. $\hat{y}$ is assigned a value of 0 if the model’s predicted probability of ’Increase’ is greater than that of ’Decrease’ and the actual polarity is ’Decrease’. The metric AUC_polarity is calculated using the Equation(6):
$AUC_{polarity}=\frac{\sum\text{sign}(p_{a}-p_{i})\times\hat{y}}{N}$ (6)
Here, $N$ is the total number of samples, and $\text{sign}(p_{a}-p_{i})$ calculates the sign of the difference in the predicted probabilities of ‘Increase’ and ‘Decrease’. This metric sensitively reflects the model’s ability to discern polar edges, even in the presence of ambiguous or weak polarity signals.
4 Conclusion
In this study, we introduced a novel deep graph model, RGCNTD (Relational Graph Convolutional Network with Tensor Decomposition), designed to predict signed interactions in biological networks. A key contribution of this work lies in its ability to handle both polar and non-polar edges, which naturally coexist in real-world biological networks. Our dataset revealed multiple types of relationships, with a significant proportion exhibiting polar characteristics, such as ‘Increase’ and ‘Decrease’, while others remained non-polar, reflecting general or undefined interactions. The coexistence of these relations poses unique challenges for predictive models, which our approach effectively addresses.
By incorporating the Conflict-Link (CL) sampling strategy, we enhanced the model’s ability to resolve polarity conflicts and improve high-confidence predictions. Compared to baseline models, RGCNTD demonstrated superior performance, particularly in distinguishing interactions with opposing polarities. The introduction of novel metrics, AUC_polarity and CP@500, was instrumental in evaluating the model’s capability to differentiate between polar and non-polar edges. These metrics provided deeper insights into the model’s predictive power within complex biological networks where diverse relationships coexist and interact.
Modeling such complexity is crucial, as it better reflects the true nature of biological systems, where multiple types of interactions occur simultaneously. This work advances our understanding of drug-gene interactions and emphasizes the importance of polarity-aware modeling in biological research. Accurately distinguishing between interaction types has significant implications for interpreting biological processes and guiding downstream applications.
This study highlights the practical significance of developing models capable of handling diverse relationships, offering a more comprehensive and accurate representation of biological networks. Future research can build on this by exploring improved strategies for resolving conflicting predictions and expanding the application of polarity-aware methods to broader biological and clinical contexts.
5 Ethics Statement
6 Data and Code Availability Statement
7 Appendix
@@ 7.0.1 Graph Convolutional Encoder
In this study, we employed a Graph Convolutional Encoder designed to capture and encode the intricate cooperative interactions between chemicals and genes. The graph convolutional encoder utilizes a hierarchical graph convolutional network architecture that iteratively aggregates and propagates information across nodes to learn node embeddings. The input to this architecture is a graph constructed from nodes represented as one-hot vectors and edges represented through adjacency matrices. The encoder opts for a two-layer GCN optimized using Chebyshev polynomials, which has proven to exhibit enhanced performance, as evidenced in previous studies. The feature update for a node $v_{i}$ is given by Equation(7):
@@
Here, $h_{i}$ represents the final node embedding, with $h_{i}^{C}$ and $h_{i}^{G}$ respectively denoting the original embeddings of chemicals and genes. We merge these two types of node embeddings through a simple vector concatenation operation $\oplus$, instead of the multi-subgraph structure mentioned in the previous model. This design choice is based on the characteristics and needs of our data, aiming to simplify the model structure while preserving key information. Our model utilizes the hierarchical structure of the graph convolutional network during the encoding phase, first calculating the embeddings for each type of node separately and then combining these embeddings through the concatenation operation. Such a structure allows the model to reduce computational complexity without sacrificing performance, enabling us to train on large-scale data.